OpenStudy (anonymous):

find y' if y=tanx/1+cosx

OpenStudy (dumbcow):

use quotient rule \[(\frac{f}{g})' = \frac{f'*g - f*g'}{g^{2}}\]

OpenStudy (anonymous):

can u show me how to apply the quotient rule plz

OpenStudy (dumbcow):

ok let \[f = \tan x\] \[f' = \sec^{2} x\] \[g = 1+\cos x\] \[g' = -\sin x\] plug it in \[= \frac{\sec^{2} x (1+\cos x) - \tan x (-\sin x)}{(1+\cos x)^{2}}\] simplify \[=\frac{\sec^{2} x + \sec x + \tan x \sin x}{(1+\cos x)^{2}}\]

OpenStudy (anonymous):

thank u

OpenStudy (dumbcow):

yw